# Performance Comparison of Gate-Based and Adiabatic Quantum Computing for Power Flow Analysis

### v1

@@ III Quantum Approximate Optimization Algorithm
The Quantum Approximate Optimization Algorithm (QAOA) belongs to the gate-based quantum computing (GQC) paradigm and is designed for solving combinatorial optimization problems. QAOA encodes problems as an Ising model or an equivalent quadratic unconstrained binary optimization (QUBO) representation, where the goal is to minimize a cost Hamiltonian defined over spin/binary variables. QAOA uses a parameterized sequence of discrete quantum gates that alternate between applying a cost Hamiltonian and a mixer Hamiltonian. Note that the cost Hamiltonian encodes the objective function of the given combinatorial optimization problem, while the mixer Hamiltonian drives transitions between computational basis states to ensure exploration of the solution space rather than getting stuck in a single configuration. By optimizing the gate parameters classically, the algorithm prepares a quantum state with a high probability of measuring near-optimal solutions. The algorithm’s depth controls the trade-off between solution quality and circuit complexity. While a higher depth can improve approximations, it also increases circuit complexity and may hinder the convergence of the classical optimizer.
@@
QAOA has been experimentally demonstrated on near-term GQC hardware, such as superconducting qubit platforms (e.g., [21]). While current devices are constrained by qubit number and gate fidelity, they are flexible in encoding arbitrary cost Hamiltonians without the embedding restrictions of adiabatic quantum computing (AQC). In principle, higher-order interactions can be incorporated into the cost Hamiltonian, though often at the expense of additional qubits or circuit depth. The performance of QAOA depends critically on both the quantum hardware quality and the classical optimizer that tunes the variational parameters to maximize solution probability.
IV Quantum Annealing
@@
Quantum annealing is a metaheuristic approach to solving combinatorial optimization problems by exploiting quantum mechanical effects. Similar to classical simulated annealing, it seeks to find the global minimum of an energy landscape; however, it uses quantum tunneling to escape local minima more efficiently. A combinatorial optimization problem is typically formulated as an Ising model or an equivalent quadratic unconstrained binary optimization (QUBO) representation. The system is composed of interacting qubits that encode the problem variables. It evolves according to a time-dependent problem Hamiltonian, starting from an initial superposition of all possible states and gradually converging to the ground state, which encodes the optimal solution. This evolution is governed by the adiabatic theorem, which ensures that slow enough changes in the problem Hamiltonian preserve the system in its ground state [20, 22].
@@
Specialized hardware implementations of quantum annealing have been developed, also known as Ising machines. D-Wave’s Advantage™ system (QA) is among the most prominent examples, featuring over 5,000 qubits and 35,000 couplers, where qubits interact according to programmable coefficients that define the problem Hamiltonian, with connectivity constraints requiring careful mapping of logical problem graphs onto the hardware topology. These limitations make embedding a critical step, as it can introduce overhead and reduce effective precision. Fujitsu’s Digital Annealer (DA) is another example. It represents an application-specific complementary metal-oxide semiconductor (CMOS) hardware555www.fujitsu.com/global/services/business-services/digital-annealer that emulates annealing behavior at room temperature. Unlike QA, which relies on cryogenic qubits, DA supports massively parallel simulated annealing and allows full connectivity between binary decision variables. Its latest generation, i.e., Quantum-Inspired Integrated Optimization software (QIIO), can handle tens of thousands of binary decision variables with high numerical precision and introduces advanced techniques, such as parallel tempering, which runs multiple replicas at different temperatures to avoid local minima [23].
@@
V Results
We propose the combinatorial PF analysis, based on which an Ising model is developed. The problem Hamiltonian is solved using the two quantum computing paradigms: (i) gate-based quantum computing (GQC), and (ii) adiabatic quantum computing (AQC). Experiments are conducted on a standard 4-bus test system consisting of one slack bus and three load buses. For a given load scenario, the combinatorial PF analysis is executed across three solvers, i.e., D-Wave’s Advantage™ system (QA), Fujitsu’s Quantum-Inspired Integrated Optimization software (QIIO), and the Quantum Approximate Optimization Algorithm (QAOA).
V-A Model Setup
For the GQC experiments, we implement QAOA using PennyLane’s lightning.qubit statevector simulator, a high-performance statevector backend that supports efficient simulation of mid-scale quantum circuits. The QAOA ansatz with $p$ alternating layers of cost and mixer Hamiltonians/unitaries is defined as:
@@ V-A Model Setup
where $H_{M}=\sum_{i=1}^{2N}X_{i}$ is the mixer Hamiltonian with Pauli-$X$ operators acting on each qubit, and $H_{C}$ is the problem-specific cost Hamiltonian defined in 5. The system size is $2N$, where $N$ is the number of buses in the PF equations 3, since both $\mu$ and $\omega$ variables are discretized. The parameters $\boldsymbol{\gamma}=(\gamma_{1},\dots,\gamma_{p})$ and $\boldsymbol{\beta}=(\beta_{1},\dots,\beta_{p})$ are randomly initialized within $0,2\pi]$, and optimized iteratively to minimize the expected energy $\langle H_{C}\rangle$. Each expectation value is estimated from projective measurements in the computational basis, with 1,000 shots per evaluation.
@@
TABLE I: QAOA implementation parameters for the combinatorial PF analysis based on the 4-bus test system.
@@
For the AQC experiments, two Ising machines are used, i.e., QA and QIIO. This implementation directly corresponds to line 10 in Algorithm 1, where the quantum/digital annealer is called. For QA, the problem Hamiltonian is mapped to the QA hardware graph using minor embedding. The chain strength is tuned to balance between preventing chain breaks and preserving the weight hierarchy of the quadratic interactions. For both approaches, 1,000 readouts/samples are collected from the annealing run, and the sample with the lowest energy is selected as the minimized solution to 5. Table II summarizes the parameters used for QA and QIIO based on the 4-bus test system. Detailed information about the quantum annealing implementation can be found in our prior studies, e.g., [8, 6, 12].
TABLE II: Quantum annealing implementation parameters for the combinatorial PF analysis based on the 4-bus test system.
@@
V-B Model Performance
Table III summarizes the computational details of solving the combinatorial PF analysis using QA, QIIO, and QAOA based on the 4-bus test system. The active and reactive power demands for load buses, $\mathbf{P}^{\mathrm{D}}$ and $\mathbf{Q}^{\mathrm{D}}$, and $\mu_{0}$ and $\omega_{0}$ for the slack bus are specified, while $\mu_{i}$ and $\omega_{i}$ are unknown for all load buses $\forall i\in\{1,2,3\}$.
The Ising model implementation results in 26 spin variables for QA, where higher-order interactions are reduced to quadratic terms using PyQUBO, and 20 decision variables for QIIO. The discrepancy in the number of variables reflects the different strategies of the underlying software frameworks in handling higher-order interactions. For QAOA, the variable count corresponds to the number of qubits. With two variables per bus $i$ (one for $\mu_{i}$ and one for $\omega_{i}$), the 4-bus test system requires 8 qubits. Compilation times for QA and QIIO are comparable, as both involve classical preprocessing and reduction of higher-order terms. QAOA, in contrast, exhibits a compile time that is one order of magnitude higher, primarily due to circuit transpilation and parameter initialization, as shown in Table III.
@@ V-B Model Performance
TABLE III: Computational detail for the combinatorial PF analysis solved with QA, QIIO, and QAOA based on the 4-bus test system.
@@
Table IV summarizes the bus complex voltages ($\mu_{i}+j\omega_{i}$) obtained from the combinatorial PF analysis using QA, QIIO, and QAOA, in comparison with the Newton–Raphson (NR) method for the 4-bus test system. The slack bus ($i=0$) with fixed values $\mu_{0}=1$ and $\omega_{0}=0$ is not shown. Both QA and QIIO reproduce the NR solution with high accuracy, with deviations on the order of $10^{-3}$ for both $\vec{\mu}$ and $\vec{\omega}$. QAOA achieves comparable accuracy for $\vec{\mu}$, but shows slightly larger deviations for $\vec{\omega}$ and does not converge to the predefined threshold within 300 iterations.
TABLE IV: Performance comparison of QA, QIIO, and QAOA with the Newton-Raphson (NR) solver for the 4-bus test system. The slack bus $i=0$ with known $\mu_{0}=1$ and $\omega_{0}=0$ is not shown.
@@
Fig. 1 shows the evolution of bus complex voltage $\mathbf{\mu}=[\mu_{1},\mu_{2},\mu_{3}]$ and $\mathbf{\omega}=[\omega_{1},\omega_{2},\omega_{3}]$ obtained with QA, QIIO, and QAOA for the 4-bus test system over iterations. The slack bus values $\mu_{0}=1$ and $\omega_{0}=0$ are not shown. Results from the NR are included as references. Fig. 1 (a–c) show the convergence of $\mu_{1-3}$, where $\mu_{2}$ reaches its NR value faster than $\mu_{1}$ and $\mu_{3}$. Fig. 1 (d–f) display the corresponding $\omega_{1-3}$, with $\omega_{2}$ converging more rapidly than $\omega_{1}$ and $\omega_{3}$. QA oscillates around the NR values for $\omega_{1-3}$ before stabilizing. Its longer overall convergence rate is mainly due to the slower convergence of $\mu_{1}$ and $\mu_{3}$. Among the methods, QIIO converges the fastest, while QAOA is the slowest. Nevertheless, all three approaches ultimately yield solutions consistent with the NR benchmark.
@@
VI Discussion
The following observations can be made:
@@ VI Discussion
  • On real quantum hardware, i.e., D-Wave’s Advantage™ system (QA), repeated executions may yield different results because of device conditions, such as noise levels, calibration, or qubit coherence, can vary over time. In addition, QA is prone to disconnection. for example, we repeatedly encountered the error “Remote end closed connection without response” when using Advantage2_system1.5 and Advantage_system6.4 for larger test systems beyond the 4-bus test system.
@@
  • For the standard 4-bus test system, we conducted experiments on both Advantage2_system1.5 and Advantage_system6.4. The former proved to be up to $20\%$ faster per iteration and consistently provided more optimal results compared to the latter.
@@
  • The choice of a small test system reflects current limitations in gate-based quantum computing (GQC) and the associated computational cost. In contrast, adiabatic quantum computing (AQC) approaches can already address larger problem sizes. For example, power systems with up to 1354 buses have been solved using Fujitsu’s Digital Annealer in our previous work [6].
VII Conclusion
In this study, we realize the combinatorial power flow (PF) analysis on both gate-based quantum computing (GQC) and adiabatic quantum computing (AQC) hardware. This reformulation yields an Ising model, which is then addressed by different solvers. For GQC, Quantum Approximate Optimization Algorithm (QAOA) is implemented using PennyLane’s lightning.qubit statevector simulator. For AQC, D-Wave’s Advantage™ system (QA) and Fujitsu’s Quantum-Inspired Integrated Optimization software (QIIO) are used. Experiments are based on a standard 4-bus test system consisting of one slack bus and three load buses. The results are evaluated in terms of solution accuracy and computational time.
@@ VII Conclusion
The results show that both QA and QIIO converge within the predefined threshold before reaching the maximum iteration limit, with QIIO achieving the fastest performance in terms of iteration count. In contrast, QAOA remains limited in accuracy under the tested configuration. Nevertheless, QA, QIIO, and QAOA all yield solutions that closely approximate the NR benchmark, confirming that the proposed combinatorial PF analysis is consistent with the classical PF equations.

### v2

@@ 3 Quantum Approximate Optimization Algorithm
QAOA belongs to the GQC paradigm and is designed for solving combinatorial optimization problems. QAOA encodes problems as an Ising model or an equivalent QUBO, and the goal is to minimize a cost Hamiltonian defined over spin/binary variables. QAOA uses a parameterized sequence of discrete quantum gates that alternate between applying a cost Hamiltonian and a mixer Hamiltonian. Note that the cost Hamiltonian encodes the objective function of the given combinatorial optimization problem, while the mixer Hamiltonian drives transitions between computational basis states to ensure exploration of the solution space rather than getting stuck in a single configuration. By optimizing the gate parameters classically, the algorithm prepares a quantum state with a high probability of yielding near-optimal solutions when measured. The algorithm’s depth controls the trade-off between solution quality and circuit complexity. While a higher depth can improve approximations, it also increases circuit complexity and may hinder the convergence of the classical optimizer.
@@
QAOA has been experimentally demonstrated on near-term GQC hardware, such as superconducting qubit platforms (e.g., Weidenfeller et al. (2022)). While current devices are constrained by the number of qubits and gate fidelity, they are flexible in encoding an arbitrary cost Hamiltonian without the embedding restrictions of AQC. In principle, higher-order interactions can be incorporated into the cost Hamiltonian, though often at the expense of additional qubits or circuit depth. The performance of QAOA depends critically on both the quality of the quantum hardware and the classical optimizer that tunes the variational parameters to maximize the solution probability.
4 Quantum Annealing
@@
Quantum annealing is a metaheuristic for solving combinatorial optimization problems by exploiting quantum-mechanical effects. Similar to classical simulated annealing, it seeks to find the global minimum of an energy landscape; however, it uses quantum tunneling to escape local minima more efficiently. A combinatorial optimization problem is typically formulated as an Ising model or an equivalent QUBO representation. The system is composed of interacting qubits that encode the problem variables. It evolves according to a time-dependent problem Hamiltonian, starting from an initial superposition of all possible states and gradually converging to the ground state, which encodes the optimal solution. This evolution is governed by the adiabatic theorem, which ensures that sufficiently slow changes in the problem Hamiltonian keep the system in its ground state Goto et al. (2019); Lucas (2014).
@@
Specialized hardware implementations of quantum annealing, also known as Ising machines, have been developed. QA is among the most prominent examples, featuring over 5,000 qubits and 35,000 couplers, in which qubits interact according to programmable coefficients that define the problem Hamiltonian, with connectivity constraints requiring careful mapping of the logical problem graph onto the hardware topology. These limitations make embedding a critical step, as it can introduce overhead and reduce effective precision. DA is another example. It represents an application-specific complementary metal-oxide semiconductor (CMOS) hardware555www.fujitsu.com/global/services/business-services/digital-annealer that emulates annealing behavior at room temperature. Unlike QA, which relies on cryogenic qubits, DA supports massively parallel simulated annealing and allows full connectivity between binary decision variables. Its latest generation, i.e., QIIO, can handle tens of thousands of binary decision variables with high numerical precision and introduces advanced techniques, such as parallel tempering, which runs multiple replicas at different temperatures to avoid local minima Yin et al. (2023).
@@
5 Results
We propose the combinatorial PF analysis, based on which an Ising model is developed. The problem Hamiltonian is solved using the two quantum computing paradigms: GQC and AQC. Experiments are conducted on a standard 4-bus test system consisting of one slack bus and three load buses. For a given load scenario, the combinatorial PF analysis is run across three solvers: QA, QIIO, and QAOA.
5.1 Model Setup
For the GQC experiments, we implement QAOA using PennyLane’s lightning.qubit statevector simulator, a high-performance statevector backend that efficiently simulates mid-scale quantum circuits. The QAOA ansatz with $p$ alternating layers of cost and mixer Hamiltonians/unitaries is defined as:
@@ 5.1 Model Setup
where $H_{M}=\sum_{i=1}^{2N}X_{i}$ is the mixer Hamiltonian with Pauli-$X$ operators acting on each qubit, and $H_{C}$ is the problem-specific cost Hamiltonian defined in ˜5. The system size is $2N$, where $N$ is the number of buses in the PF equations ˜3, since both $\mu$ and $\omega$ variables are discretized. The parameters $\boldsymbol{\gamma}=(\gamma_{1},\dots,\gamma_{p})$ and $\boldsymbol{\beta}=(\beta_{1},\dots,\beta_{p})$ are randomly initialized within $0,2\pi]$, and optimized iteratively to minimize the expected energy $\langle H_{C}\rangle$. Each expectation value is estimated from projective measurements in the computational basis, with 1,000 shots per evaluation.
@@
Table 1: QAOA implementation parameters for the combinatorial PF analysis based on the 4-bus test system.
@@
For the AQC experiments, two Ising machines are used, i.e., QA and QIIO. This implementation directly corresponds to line 10 in Algorithm˜1, where the quantum/digital annealer is called. For QA, the problem Hamiltonian is mapped to the QA hardware graph using minor embedding. The chain strength is tuned to balance between preventing chain breaks and preserving the weight hierarchy of the quadratic interactions. For both approaches, 1,000 readouts/samples are collected from the annealing run, and the sample with the lowest energy is selected as the minimized solution to ˜5. Table 2 summarizes the parameters used for QA and QIIO based on the 4-bus test system. Detailed information about the quantum annealing implementation can be found in our prior studies, e.g., Kaseb et al. (2024b, 2025a, 2025b).
Table 2: Quantum annealing implementation parameters for the combinatorial PF analysis based on the 4-bus test system.
@@
5.2 Model Performance
Table˜3 summarizes the computational details of solving the combinatorial PF analysis using QA, QIIO, and QAOA based on the 4-bus test system. The active and reactive power demands for load buses, $\mathbf{P}^{\mathrm{D}}$ and $\mathbf{Q}^{\mathrm{D}}$, and $\mu_{0}$ and $\omega_{0}$ for the slack bus are specified, while $\mu_{i}$ and $\omega_{i}$ are unknown for all load buses $\forall i\in\{1,2,3\}$.
The Ising model implementation results in 26 spin variables for QA, where higher-order interactions are reduced to quadratic terms using PyQUBO, and 20 decision variables for QIIO. The discrepancy in the number of variables reflects the different strategies of the underlying software frameworks in handling higher-order interactions. For QAOA, the variable count corresponds to the number of qubits. With two variables per bus $i$ (one for $\mu_{i}$ and one for $\omega_{i}$), the 4-bus test system requires 8 qubits. Compilation times for QA and QIIO are comparable, as both involve classical preprocessing and reduction of higher-order terms. QAOA, in contrast, exhibits a compile time that is one order of magnitude higher, primarily due to circuit transpilation and parameter initialization, as shown in Table˜3.
@@ 5.2 Model Performance
Table 3: Computational detail for the combinatorial PF analysis solved with QA, QIIO, and QAOA based on the 4-bus test system.
@@
Table˜4 summarizes the bus complex voltages ($\mu_{i}+j\omega_{i}$) obtained from the combinatorial PF analysis using QA, QIIO, and QAOA, in comparison with the NR method for the 4-bus test system. The slack bus ($i=0$) with fixed values $\mu_{0}=1$ and $\omega_{0}=0$ is not shown. Both QA and QIIO reproduce the NR solution with high accuracy, with deviations on the order of $10^{-3}$ for both $\vec{\mu}$ and $\vec{\omega}$. QAOA achieves comparable accuracy for $\vec{\mu}$, but shows slightly larger deviations for $\vec{\omega}$ and does not converge to the predefined threshold within 300 iterations.
Table 4: Performance comparison of QA, QIIO, and QAOA with the NR solver for the 4-bus test system. The slack bus $i=0$ with known $\mu_{0}=1$ and $\omega_{0}=0$ is not shown.
@@
Fig.˜2 shows the evolution of bus complex voltage $\mathbf{\mu}=[\mu_{1},\mu_{2},\mu_{3}]$ and $\mathbf{\omega}=[\omega_{1},\omega_{2},\omega_{3}]$ obtained with QA, QIIO, and QAOA for the 4-bus test system over iterations. The slack bus values $\mu_{0}=1$ and $\omega_{0}=0$ are not shown. Results from the NR are included as references. Fig.˜2 (a–c) show the convergence of $\mu_{1-3}$, where $\mu_{2}$ reaches its NR value faster than $\mu_{1}$ and $\mu_{3}$. Fig.˜2 (d–f) display the corresponding $\omega_{1-3}$, with $\omega_{2}$ converging more rapidly than $\omega_{1}$ and $\omega_{3}$. QA oscillates around the NR values for $\omega_{1-3}$ before stabilizing. Its longer overall convergence rate is mainly due to the slower convergence of $\mu_{1}$ and $\mu_{3}$. Among the methods, QIIO converges the fastest, while QAOA is the slowest. Nevertheless, all three approaches ultimately yield solutions consistent with the NR benchmark.
@@
6 Discussion
Quantum computing is an emerging field in which both hardware and software are undergoing rapid development. Although recent advances in quantum hardware have progressed faster than anticipated only a few years ago, current devices are not yet capable of supporting large-scale and practical applications. Nevertheless, without well-suited algorithms and software frameworks, even fully developed quantum processors would have limited practical value. Consequently, research on quantum algorithms for application domains where computational performance is critical must progress in parallel with hardware development. At present, despite the availability of early quantum hardware, electricity grid operations cannot meaningfully benefit from it because quantum algorithms tailored to power system problems are still in development. Motivated by this gap, this paper investigates two major quantum computing paradigms, GQC and AQC, for solving the PF problem. Based on the results presented in this work, the following observations can be made:
@@ 6 Discussion
  • On real quantum hardware, i.e., QA, repeated executions may yield different results because device conditions, such as noise levels, calibration, or qubit coherence, can vary over time. In addition, QA is prone to disconnection. For example, we repeatedly encountered the error “Remote end closed connection without response” when using Advantage2_system1.5 and Advantage_system6.4 for larger test systems beyond the 4-bus test system.
@@
  • For the standard 4-bus test system, we conducted experiments on both Advantage2_system1.5 and Advantage _system6.4. The former proved to be up to $20\%$ faster per iteration and consistently produced better results than the latter.
@@
  • The choice of a small test system reflects current limitations in GQC and the associated computational cost. In contrast, AQC approaches can already address larger problem sizes. For example, power systems with up to 1354 buses have been solved using DA in our previous work Kaseb et al. (2025a).
7 Conclusion
@@
In this study, we implement combinatorial PF analysis on both GQC and AQC hardware. The PF problem is reformulated as an Ising model, which is then solved using different quantum solvers. For GQC, QAOA is implemented using PennyLane’s lightning.qubit statevector simulator. For AQC, both QA and QIIO are employed. Experiments are conducted on a standard 4-bus test system consisting of one slack bus and three load buses. The results are evaluated in terms of solution accuracy and computational performance, and indicate that both QA and QIIO converge within the predefined threshold before reaching the maximum iteration limit, with QIIO achieving the fastest performance in terms of iteration count. In contrast, QAOA exhibits limited accuracy under the tested configuration. However, QA, QIIO, and QAOA all produce solutions that closely approximate the NR benchmark, confirming that the proposed combinatorial PF formulation is consistent with the classical AC PF equations.
Acknowledgment
This work is part of the DATALESs project (project no. 482.20.602), jointly financed by the Netherlands Organization for Scientific Research (NWO) and the National Natural Science Foundation of China (NSFC). The authors gratefully acknowledge Heinz Wilkening and the service contract "Quantum Computing for Load Flow" (contract no. 690523) with the European Commission Directorate-General Joint Research Centre (EC DG JRC). The authors also like to thank Fujitsu Technology Solutions for providing access to the QIIO software [^6] https://en-portal.research.global.fujitsu.com/kozuchi. Furthermore, the authors acknowledge TNO for the access to TNO’s Quantum Application Lab Facility, with special thanks to Frank Phillipson for his assistance.
[…]